# On robustness of Spectral Rényi divergence

### v3

@@ Appendix B Proof of Theorem 3
=$=\left\{\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\theta^{(0)}\right\}-\left\{\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]-\theta^{(0)}\right\}$
@@
$\left\|\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]\right\|$ =$=\Bigg{\|}\left\{\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]\right\}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}])$
@@
$\quad+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
@@
$\quad+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)$
$=\frac{\alpha}{(1-\alpha)n}(T_{1}+T_{2})+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\Theta}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
@@
$\leq\frac{1}{(1-\alpha)n}\sup_{\theta,\omega}\|\nabla_{\theta}\log S_{\theta}(\omega)\|$
$\,\,+\varepsilon\Bigg{\{}\frac{1}{(1-\alpha)}\Big{(}\sup_{\theta,\omega}\|\nabla_{\theta,\omega}S_{\theta}(\omega)\|\Big{)}^{2}+\frac{2}{(1-\alpha)}\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\Bigg{\}}$
@@
$\leq\frac{\gamma_{m}}{(1-\alpha)n}\sup_{\theta,\omega}\|\nabla_{\theta}\log S_{\theta}(\omega)\|$
$\qquad+\varepsilon\Bigg{\{}1+\frac{\gamma_{m}}{(1-\alpha)}\Big{(}\sup_{\theta,\omega}\|\nabla_{\theta,\omega}S_{\theta}(\omega)\|\Big{)}^{2}+\frac{2\gamma_{m}}{(1-\alpha)}\sup_{\theta\in\Theta}\frac{1}{n}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\Bigg{\}}.$
@@ Appendix C Proof of Proposition 2
We begin with utilizing the descent lemma (c.f., [2]):
@@
[…]
@@ Appendix D Proof of Proposition 3
Proof.
The difference between $\widehat{\theta}^{(1)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]$ and $\widehat{\theta}^{(1)}_{1}[\widetilde{I}_{n}]$ is explicitly written as
$\|\widehat{\theta}^{(1)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{1}[\widetilde{I}_{n}]\|$ =$=\|\{\widehat{\theta}^{(1)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]-\theta^{(0)}\}-\{\widehat{\theta}^{(1)}_{1}[\widetilde{I}_{n}]-\theta^{(0)}\}\|$
@@
$\sum_{\omega\in\Omega_{n}}\left[1-\frac{\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}{S_{\theta}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega)\Bigg{|}_{\theta=\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}$
$=\sum_{\omega\in\Omega_{n}}\left[1-\frac{\widetilde{I}_{n}(\omega)}{S_{\theta}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega)\Bigg{|}_{\theta=\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}-z\frac{\nabla_{\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}\log S_{\theta}(\omega^{*})}{S_{\widehat{\theta}^{(\infty)}_{1}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega^{*})}.$
@@
Appendix E Additional simulation studies
@@
Tables 3 and 4 show the estimation results for the Brune spectral model with attenuation in Section 4.1 on the basis of the different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$. For almost all the settings of initial values, the spectral density based on the spectral $\alpha=0.5$-Rényi divergence performs the best.
@@
IS with $I_{n}$ $\theta^{(0),4}$
IS with $I_{n}^{S}$ $\theta^{(0),4}$
R ($\alpha=0.50$) $\theta^{(0),5}$0.01($\pm$0.06) -0.18 ($\pm$0.05) 0.47 ($\pm$0.16)
@@
IS with $I_{n}^{S}$ $\theta^{(0),5}$0.28 ($\pm$0.16) 1.84 ($\pm$1.66) 1.90 ($\pm$0.63)
Table 4: The mean values of biases with standard deviations with the trigonometric trends. For each initial value, the values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init.
@@
IS with $I_{n}$ $\theta^{(0),4}$
IS with $I_{n}^{S}$ $\theta^{(0),4}$
@@
Figures 4 and 5 display an instance of a set of the estimated spectral densities. In each figure, the following representations are used:
@@
  • the solid blue curve illustrates the spectral density with the minimum Itakura–Saito divergence estimate plugged-in;
  • the solid red, salmon pink, and green curves display the spectral density with the minimum spectral Rényi divergence ($\alpha=0.9,0.75,0.5$) estimates plugged-in, respectively.
@@
Figure 4: Spectral densities with the estimates plugged-in for the Brune spectral model with attenuation and with the trigonometric trends of $z_{1}=z_{2}=2.5$. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. The red dashed lines denote the frequencies at which the outliers are injected. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Figure 5: Spectral densities with the estimates plugged-in for the Brune spectral model with attenuation and with the trigonometric trends of $z_{1}=z_{2}=25$. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. The red dashed lines denote the frequencies at which the outliers are injected. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.

### v4

@@ Appendix B Proof of Theorem 3
$=\left\{\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\theta^{(0)}\right\}-\left\{\widehat{\theta}^{(1)}_{\alpha}[\widetilde{I}_{n}]-\theta^{(0)}\right\}$
@@
$\left\|\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}]\right\|$ $=\Bigg\|\left\{\widehat{\theta}^{(m+1)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}^{z,\omega^{*}}]\right\}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad(\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(m)}_{\alpha}[\widetilde{I}_{n}])$
@@
$\quad+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\mathrm{int}(\Theta)}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
@@
$\quad+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\mathrm{int}(\Theta)}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right)$
$=\frac{\alpha}{(1-\alpha)n}(T_{1}+T_{2})+O\left(\frac{2\varepsilon}{(1-\alpha)n}\sup_{\theta\in\mathrm{int}(\Theta)}\sum_{\omega\in\Omega_{n}}\|\nabla^{2}_{\theta}\log S_{\theta}(\omega)\|_{\mathrm{op}}\right),$
@@
$=\frac{1}{(1-\alpha)n}O(U_{1})+\varepsilon\Bigg\{\frac{1}{(1-\alpha)}O(U_{1}^{2})+\frac{2}{(1-\alpha)}O(U_{2})\Bigg\}$
@@
$\leq\frac{\gamma_{m}}{(1-\alpha)n}O(U_{1})+\varepsilon\Bigg\{1+\frac{\gamma_{m}}{(1-\alpha)}O(U_{1}^{2})+\frac{2\gamma_{m}}{(1-\alpha)}O(U_{2})\Bigg\}.$
@@ Appendix C Proof of Proposition 2
We begin with utilizing the descent lemma (c.f., Bertsekas (1999)):
@@
where $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner-product. From equation (12), we have
@@ Appendix D Proof of Proposition 3
The difference between $\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]$ and $\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}_{n}]$ is explicitly written as
$\|\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]-\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}_{n}]\|$ $=\|\{\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]-\theta^{(0)}\}-\{\widehat{\theta}^{(1)}_{\mathrm{IS}}[\widetilde{I}_{n}]-\theta^{(0)}\}\|$
@@
$\sum_{\omega\in\Omega_{n}}\left[1-\frac{\widetilde{I}^{z,\omega^{*}}_{n}(\omega)}{S_{\theta}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega)\Bigg|_{\theta=\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}$
$=\sum_{\omega\in\Omega_{n}}\left[1-\frac{\widetilde{I}_{n}(\omega)}{S_{\theta}(\omega)}\right]\nabla_{\theta}\log S_{\theta}(\omega)\Bigg|_{\theta=\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}-z\frac{\nabla_{\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}\log S_{\theta}(\omega^{*})}{S_{\widehat{\theta}^{(\infty)}_{\mathrm{IS}}[\widetilde{I}^{z,\omega^{*}}_{n}]}(\omega^{*})}.$
@@
Appendix E Additional comparison of optimization paths
This appendix provides an additional comparison of optimization paths. Here we consider contaminated observations defined as
$X_{t}^{\circ}=X_{t}+\sqrt{z}\sin(t\pi/2),\quad t=1,\ldots,n=500$
with $(X_{t})$ generated from AR(2) model having $(\sigma,\varphi_{1},\varphi_{2})$.
Figure 7: Comparison of optimization paths using the AR(2) model with the length $n$ of a time series set to $500$. Panels (a) and (c) show the optimization paths (curves with light colors) of the spectral Rényi divergence minimization that start from the initial value (the gray circle) and terminate at points after $10000$ iterations (the inverted triangles), where the colors indicate different values of the contamination $z$. Panels (b) and (d) show the corresponding results for the Itakura–Saito divergence minimization.
To remove the parameter constraints arising from the stationarity condition, we employ the following transformation:
$\theta_{2}:=\,\mathrm{atanh}\left(\frac{\varphi_{1}}{1-\varphi_{2}}\right)\in\mathbb{R}\quad\text{and}\quad\theta_{3}:=\,\mathrm{atanh}(\varphi_{2})\in\mathbb{R}.$ (16)
This transformation is interpreted as follows. First, we transform the AR coefficients $\varphi_{1},\varphi_{2}$ into the partial autocorrelation (PARCOR) coefficients $\kappa_{1},\kappa_{2}$ using the Durbin–Levinson formula:
$\kappa_{1}:=\frac{\varphi_{1}}{1-\varphi_{2}}\in(-1,1)\quad\text{and}\quad\kappa_{2}:=\varphi_{2}\in(-1,1).$
Next, we map the PARCOR coefficients to unrestricted real values via the Fisher transformation:
$\theta_{2}:=\,\mathrm{atanh}\left(\kappa_{1}\right)\in\mathbb{R}\quad\text{and}\quad\theta_{3}:=\,\mathrm{atanh}(\kappa_{2})\in\mathbb{R},$
which yields the transformation in (16). We perform 10000 iterations of gradient descent with a step size of 0.01 with respect to $(\log\sigma,\theta_{2},\theta_{3})$.
Figure 7 shows a comparison of the optimization paths for the AR(2) model, where the true parameter values are set to $(\sigma^{*},\varphi^{*}_{1},\varphi^{*}_{2})=(1,0.9,-0.2)$ and the hyperparameter $\alpha$ is set to $\alpha=0.9$. Even for the PARCOR coefficients (in addition to $\sigma$), the optimization trajectory based on the Itakura–Saito divergence remains sensitive to outliers, whereas that based on the spectral Rényi divergence remains stable with respect to all parameters.
Appendix F Additional simulation studies
@@
Tables 5 and 6 show the estimation results for the Brune spectral model with attenuation in Section 4.2 on the basis of the different initial values $\theta^{(0),4}=(0.1,0.1,0.1)$ and $\theta^{(0),5}=(2,2,2)$. For almost all the settings of initial values, the spectral density based on the spectral $\alpha=0.5$-Rényi divergence performs the best.
@@
IS with $I_{n}$ $\theta^{(0),4}$ * * *
IS with $I_{n}^{S}$ $\theta^{(0),4}$ * * *
R ($\alpha=0.50$) $\theta^{(0),5}$ -0.01($\pm$0.06) -0.18 ($\pm$0.05) 0.47 ($\pm$0.16)
@@
IS with $I_{n}^{S}$ $\theta^{(0),5}$ -0.28 ($\pm$0.16) 1.84 ($\pm$1.66) 1.90 ($\pm$0.63)
Table 6: The mean values of biases with standard deviations with the trigonometric trends. For each initial value, the values closest to zero are underlined. Rényi is abbreviated as R; Itakura–Saito is abbreviated as IS; Initial value is abbreviated as Init. Values greater than 10 are denoted by *.
@@
IS with $I_{n}$ $\theta^{(0),4}$ * * *
IS with $I_{n}^{S}$ $\theta^{(0),4}$ * * *
@@
Figures 8 and 9 display an instance of a set of the estimated spectral densities. In each figure, the following representations are used:
@@
  • the solid blue curve illustrates the spectral density with the minimum Itakura–Saito divergence estimate plugged in;
  • the solid red, salmon pink, and green curves display the spectral density with the minimum spectral Rényi divergence ($\alpha=0.9,0.75,0.5$) estimates plugged in, respectively.
@@
Figure 8: Spectral densities with the estimates plugged in for the Brune spectral model with attenuation and with the trigonometric trends of $z_{1}=z_{2}=2.5$. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. The red dashed lines denote the frequencies at which the outliers are injected. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.
Figure 9: Spectral densities with the estimates plugged in for the Brune spectral model with attenuation and with the trigonometric trends of $z_{1}=z_{2}=25$. The gray curve is the periodogram. The true spectral density is colored in black. Spectral densities based on the spectral Rényi divergence ($\alpha=0.5,0.75,0.9$) are colored in red, salmon pink, and green,respectively. The spectral density based on the Itakura–Saito divergence is colored in blue. The red dashed lines denote the frequencies at which the outliers are injected. (a) the result based on the initial value $\theta^{(0),1}$, (b) the result based on the initial value $\theta^{(0),2}$, (c) the result based on the initial value $\theta^{(0),3}$.